# SDDBench: A Benchmark for Synthesizable Drug Design

### v1

@@ 3.2.2 Evaluation of the Reliability of our Round-trip Score
In this section, we assess the reliability of . As previously discussed, a reliable should demonstrate a significant gap between molecules for which the retrosynthetic planner can predict feasible routes and those for which it cannot.
We begin by constructing a reaction network using all available data from USPTO-full. From this network, we extract synthetic routes where the leaf nodes are starting materials. This process yields approximately 105,000 reactions across all synthetic routes. We then partition these routes into training, validation, and test sets, allocating 98%, 1%, and 1% respectively, based on the target molecules. Subsequently, we train Neuralsym (Segler & Waller, 2017) using the reaction data from the training set. For a detailed description of all datasets utilized in this study, please refer to Appendix A.2.
We employ Neuralsym+Beam Search (beam size is 2) as our retrosynthetic planner to predict synthetic routes for target molecules in the test dataset. We then compare these predicted routes to the reference routes provided in the dataset. If the starting materials in our predicted routes match those in the reference routes, we categorize these molecules as successful cases. Conversely, if they don’t match, we classify them as failed cases. This approach allows us to distinguish between molecules for which our retrosynthetic planner can successfully predict feasible synthetic routes and those for which it cannot.
Next, we use a reaction model to reproduce the synthetic routes and calculate for each molecule. We then compare the values between the successful and failed cases. If is a reliable metric, we expect to see a significant difference in its values between the two groups. Specifically, a large gap in between successful and failed cases would indicate that our metric effectively distinguishes between molecules with accurately predicted synthetic routes and those without.
Our reaction model is trained on an extensive dataset of approximately 811,000 reactions. We utilize a Transformer-Decoder (Vaswani et al., 2017) architecture incorporating KV cache to enhance inference speed. The model’s input consists of a sequence of chemical reactants, from which it predicts the corresponding sequence of products. This approach aligns with previous works such as the Molecular Transformer (Schwaller et al., 2019) and Chemformer (Irwin et al., 2022). During the inference process, we prioritize accuracy by directly using the product from the corresponding reactions in the dataset when the reactants are present, rather than relying on the model’s prediction. This strategy ensures the reaction model’s reliability. Importantly, the dataset used to train our retrosynthesis model is a subset of the larger dataset used for the reaction model. This relationship guarantees that the retrosynthesis model’s predictions fall within the data distribution of the reaction model’s training set. Generally, our reaction model outperforms the retrosynthesis model in terms of prediction accuracy. The reaction model demonstrates exceptional performance, achieving a top-1 test accuracy of 90% on the USPTO-480k dataset. In contrast, the retrosynthesis model achieves a top-1 test accuracy of 50% on the USPTO-50k dataset. This difference in accuracy can be attributed to the nature of their respective tasks: reaction prediction is a discriminative task, while retrosynthesis prediction is a generation task. In the latter, the correct answer may not always be present in the reference reactions of the test dataset, making it inherently more challenging.
@@ Results.
In addition to our initial definition of successful and failed cases, we introduce a second scenario inspired by the round-trip accuracy commonly used in single-step retrosynthesis evaluation. In this scenario, we classify molecules as successful cases when and are identical, indicating that the retrosynthetic planner can predict feasible synthetic routes for these molecules. This approach doesn’t rely on the limited reference routes in the test dataset but instead depends on the accuracy of our reaction model, which we’ve previously discussed.
Figure 5 illustrates the round-trip scores for successful and failed cases under both definitions. When using the set-wise match metric as our evaluation criterion, we observe some little overlap in round-trip scores between successful and failed cases, though a significant gap remains. However, when employing the molecule-wise (round-trip) match metric, we find almost no overlap in round-trip scores between successful and failed cases, and the gap becomes even more pronounced.
Based on these results, we conclude that the proposed round-trip score is a reliable measure for distinguishing which molecules are likely to have feasible synthetic routes predicted by our retrosynthetic planner, given its training on current reaction data. Consequently, in our subsequent experimental section, we use this round-trip score to evaluate molecules proposed by Structure-Based Drug Design (SBDD) generative models. This evaluation helps us assess the likelihood that these proposed molecules can have feasible synthetic routes predicted by the retrosynthetic planner.
Ablation Study.
We conduct an evaluation of the Synthetic Accessibility (SA) Score for both successful and failed cases. As shown in Figure 6, there is a substantial overlap in SA Scores between these two groups. This significant overlap suggests that the SA Score is not an effective metric for distinguishing between molecules for which our retrosynthetic planner can successfully predict feasible routes and those for which it cannot.
Given this limitation of the SA Score, we conclude that our proposed round-trip score offers a more reliable measurement of synthetic accessibility. The round-trip score’s ability to clearly differentiate between successful and failed cases makes it a more robust and informative metric for assessing the likelihood of successful retrosynthetic planning.
@@ 4 Related Work
Structured-Based Drug Design.
Generative models for Structure-Based Drug Design (SBDD) can be broadly categorized into two main types: non-diffusion and diffusion-based models. Non-diffusion models encompass a range of approaches, including LiGAN (Ragoza et al., 2022), AR (Luo et al., 2021), Pocket2Mol (Peng et al., 2022), GraphBP (Liu et al., 2022b), FLAG (Zhang et al., 2022), and DrugGPS (Zhang & Liu, 2023). On the other hand, diffusion-based models can be considered as a alternative, such as TargetDiff (Guan et al., 2023a), DiffSBDD (Schneuing et al., 2022), and DecompDiff (Guan et al., 2023b).
@@
Reaction prediction models can be broadly categorized into two approaches: template-based and template-free. Template-based methods (Wei et al., 2016; Segler & Waller, 2017; Qian et al., 2020; Chen & Jung, 2022) begin by extracting reaction templates from a reaction database. These methods then predict the most suitable template class based on the given reactants and apply the predefined, encoded template to generate the product. Template-free approaches, on the other hand, are more diverse. Some, inspired by chemical reaction mechanisms, adopt a two-stage learning process (Jin et al., 2017). They first identify the chemical reaction centers of the reactants using atom mapping numbers, and then form new bonds or break existing ones between atoms at these centers. However, most contemporary strategies employ an end-to-end, template-free learning paradigm for reaction prediction. Several methods (Yang et al., 2019; Schwaller et al., 2019; Tetko et al., 2020; Irwin et al., 2022; Lu & Zhang, 2022; Zhao et al., 2022; Tu & Coley, 2022) frame reaction prediction as a sequence-to-sequence or graph-to-sequence problem. Other approaches (Bradshaw et al., 2019a; Do et al., 2019; Sacha et al., 2021; Bi et al., 2021; Meng et al., 2023) predict the product by directly performing graph transformations on the reactants’ graph representations.
@@
Existing retrosynthesis models (Segler & Waller, 2017; Coley et al., 2017; Liu et al., 2017; Zheng et al., 2019; Chen et al., 2019; Dai et al., 2019; Karpov et al., 2019; Chen et al., 2020; Ishiguro et al., 2020; Guo et al., 2020; Tetko et al., 2020; Shi et al., 2020; Yan et al., 2020; Seo et al., 2021; Chen & Jung, 2021; Lee et al., 2021; Seidl et al., 2021; Somnath et al., 2021; Sun et al., 2021; Yan et al., 2022; Fang et al., 2022; Gao et al., 2022b; Wan et al., 2022; He et al., 2022; Liu et al., 2022a; Tu & Coley, 2022; Lin et al., 2022; Zhong et al., 2022; Baker et al., 2023; Zhong et al., 2023; Yu et al., 2023; Li et al., 2023b; Xie et al., 2023; Sacha et al., 2023; Zhu et al., 2023b; Jiang et al., 2023; Qian et al., 2023; Xiong et al., 2023; Wang et al., 2023; Gao et al., 2023; Zhu et al., 2023a; Lan et al., 2023; Chen et al., 2023; Yao et al., 2023; Lin et al., 2023; Liu et al., 2024b; Zhang et al., 2024b, a; Lan et al., 2024) can be broadly categorized into three main types: template-free, semi-template-based, and template-based methods. These categories can be further refined based on their utilization of atom mapping information. Template-free methods typically approach retrosynthesis as either a translation problem (Karpov et al., 2019) or a graph edit problem (Sacha et al., 2021). Some of these methods optionally use atom mapping to align input and output molecules (Seo et al., 2021; Zhong et al., 2022; Yao et al., 2023). Template-based methods (Segler & Waller, 2017; Dai et al., 2019) leverage atom mapping information to create a pool of reaction templates. They generally frame retrosynthesis as a template classification or retrieval problem. Semi-template-based methods (Shi et al., 2020; Somnath et al., 2021) also utilize atom mapping information, but they employ it to obtain other prior information, such as identifying reaction centers. Many of these methods adopt a two-stage learning paradigm for retrosynthesis: First, they identify the reaction center in the product molecule and break it into synthons. Then, they transform these synthons into reactants.
@@
A variety of search algorithms have been developed to navigate the synthetic planning. These include beam search, neural A* search (Chen et al., 2020; Han et al., 2022; Xie et al., 2022), Monte Carlo Tree Search (MCTS) (Segler et al., 2018; Hong et al., 2021), and reinforcement learning-based (RL-based) search (Yu et al., 2022). Some other works to this field include (Kishimoto et al., 2019; Heifets & Jurisica, 2012; Kim et al., 2021; Hassen et al., 2022; Li et al., 2023a; Zhang et al., 2023; Liu et al., 2023a; Lee et al., 2023; Yuan et al., 2024; Tripp et al., 2024). These algorithms are designed to explore the vast reaction space, prioritizing the most promising synthetic routes.
5 Experiments
In this section, we present a comprehensive evaluation of molecules generated by Structure-Based Drug Design (SBDD) models. Our assessment primarily focuses on the aspect of synthesizability, a crucial factor in drug discovery. We analyze these molecules using two key metrics: their round-trip scores and search success rates. This approach allows us to thoroughly assess the practical feasibility of synthesizing these computer-generated molecular structures.
Baselines and Settings.
For our evaluation, we employ a diverse set of state-of-the-art SBDD models, including LiGAN (Ragoza et al., 2022), AR (Luo et al., 2021), Pocket2Mol (Peng et al., 2022), FLAG (Zhang et al., 2022), TargetDiff (Guan et al., 2023a), DrugGPS (Zhang & Liu, 2023), and DecompDiff (Guan et al., 2023b). These models are trained and tested using the CrossDocked dataset (Francoeur et al., 2020), which comprises an extensive collection of 22.5 million protein-molecule structures. Our experimental setup involves randomly selecting 100,000 protein-ligand pairs from this dataset for training purposes. For testing, we draw 100 proteins from the remaining data points. To ensure a comprehensive evaluation, we randomly sample 100 molecules for each protein pocket in the test dataset, resulting in a total of 10,000 molecules. In our evaluation, we adopt the FusionRetro (Liu et al., 2023b) setup, employing a beam size of 5 for all models to ensure a fair comparison. Our experiments are conducted using an Nvidia H100 80G GPU. We evaluate the synthesizability of 10,000 molecules, which takes at least 12 hours to complete. We find that the primary bottleneck in time complexity is Neuralsym’s retrosynthesis prediction, which requires 0.157s per prediction. In contrast, our forward model, which utilizes KV cache and batch decoding, achieves a rapid prediction time of only 0.0055s per reaction. Due to computational constraints, we are unable to use a beam size of 50 as employed in retrosynthesis evaluation (Dai et al., 2019). Besides, our approach generates about 5 synthetic routes per molecule, in contrast to previous methods in retrosynthetic planning that typically produce only one route. This offers a more comprehensive evaluation compared to the search success rate metric used in earlier studies (Chen et al., 2020) for evaluating search algorithms.
Metrics.
We assess the performance of each model using two key metrics. Top-k Route Quality: We calculate the percentage of molecules for which at least one of the top-k predicted synthetic routes achieves a round-trip score exceeding 0.9. This metric helps us gauge the quality and feasibility of the proposed synthetic routes. Search Success Rate: This metric measures the proportion of molecules for which the model successfully identifies viable synthetic routes within the specified constraints.
Results.
| Model | Search Success Rate | Top-1 Max | Top-2 Max | Top-3 Max | Top-4 Max | Top-5 Max |
| LiGAN | 9.68 | 1.30 | 1.64 | 1.84 | 1.95 | 2.06 |
| TargetDiff | 13.16 | 1.75 | 2.25 | 2.50 | 2.68 | 2.78 |
| DecompDiff | 17.26 | 2.66 | 3.30 | 3.64 | 3.89 | 3.96 |
| DrugGPS | 21.29 | 4.24 | 5.40 | 6.06 | 6.42 | 6.72 |
| AR | 27.42 | 5.13 | 6.32 | 7.01 | 7.28 | 7.56 |
| FLAG | 32.56 | 7.59 | 9.65 | 10.73 | 11.38 | 11.92 |
| Pocket2Mol | 45.98 | 11.35 | 14.55 | 16.29 | 17.28 | 18.14 |
Based on the results presented in Table 1, we can draw several conclusions. There is a significant variation in performance across different SBDD models. The Top-5 Max ranges from 2.06% for LiGAN to 18.14% for Pocket2Mol, indicating a substantial difference in the models’ abilities to generate synthetically accessible molecules. Pocket2Mol consistently outperforms other models across all metrics, with 18.14% of its generated molecules having at least one high-quality synthetic route (round-trip score ) among the top 5 predictions. The improvement in performance from Top1 to Top5 suggests that considering multiple top predictions can significantly increase the likelihood of finding feasible synthetic routes.
Notably, the performance ranking of models remains consistent across all top-k evaluations. The performance increase from Top-4 to Top-5 is less than 1%, indicating that the performance is approaching saturation. Additionally, as shown in Table 5 in Appendix B, the performance gap between models generally widens as increases, with most gaps showing an upward trend. These observations suggest that our chosen beam size of 5 is sufficient to provide an accurate ranking of each model’s performance. This consistency in ranking and the approaching saturation point lend credibility to our evaluation methodology and the reliability of our comparative analysis. More experiment results can be found in Appendix B. The analysis of molecular properties from various generative models, as presented in Table 8, reveals that superior molecular properties do not always correlate with better synthesizability. Even for the best-performing model, a considerable portion of generated molecules still lack high-quality synthetic routes, indicating room for improvement in generating synthetically accessible molecules in SBDD tasks. These findings underscore the importance of evaluating synthetic accessibility in SBDD models and highlight the potential of using top-k predictions to identify feasible synthetic routes for generated molecules.

### v2

@@ 4.1 Evaluating the Reliability of Synthesizability Metrics
We employ the template-based Neuralsym as our retrosynthesis model, training it on reactions derived from the 105,218 data points. For predicting synthetic routes for new molecules, we leverage Neuralsym (Segler & Waller, 2017) integrated with beam search as our retrosynthetic planner. We use the Transformer (Vaswani et al., 2017) Decoder as our forward reaction prediction model, training it on about 916k reactions. All experiments in this paper are conducted using an Nvidia H100 80G GPU. We find that the primary bottleneck in time complexity during the search is Neuralsym’s retrosynthesis prediction, which requires 0.157s per prediction. In contrast, our forward model, which utilizes key-value (KV) cache and batch decoding, achieves a rapid prediction time of only 0.0055s per reaction.
Evaluation Protocol.
We use 1,068 data points from the test set to evaluate the ability of the synthesizability evaluation metric to distinguish between feasible and infeasible routes. Please note that the number of molecules evaluated far exceeds the number of molecules used in the benchmarks for assessing retrosynthetic planning search algorithms (Chen et al., 2020). Our evaluation includes more than five times their number (189 testing molecules). Therefore, we believe the scale of our evaluation is highly convincing. For these 1,068 target molecules, we first employ the retrosynthetic planner to predict synthetic routes with a beam size of 5. During the search process, the depth of each route is limited to a maximum of 15, which is the highest depth of all routes in our dataset. While the planner can generate up to five different routes for each molecule, we only consider the route with the highest confidence score. As a result, our retrosynthetic planner successfully generates routes for 1,027 molecules, while failing to generate routes for 41 molecules.
To determine the feasibility of a predicted route, we compare it against the reference routes in the test set. If the starting materials of the predicted route match the starting materials of any reference route, the route is deemed feasible. However, it is important to note that the reference routes in the test set do not cover all possible feasible routes. For predicted routes that do not match any reference routes, we manually evaluate their feasibility using the Chemical Abstracts Service (CAS) SciFinder (Gabrielson, 2018) tool222https://scifinder-n.cas.org/, complemented by our domain expertise, as two of our authors are Ph.D. students focusing on organic chemistry research.
Using CAS SciFinder to evaluate route feasibility, we find that all reactions in some routes are documented in the CAS database, as illustrated in Figure 4. Furthermore, although some reactions are not explicitly listed in the CAS database, we find that they partially match reactions documented in the database, as shown in Figure 5. Therefore, based on our knowledge, we think these reactions are also accurate.
Through this process, we find that for 526 molecules, the predicted routes are identified as feasible based on the reference routes in the test set. 510 of these molecules have round-trip scores of 1. For an additional 501 molecules, feasibility is confirmed through manual evaluation and the use of CAS tools. Among these, the reactions in the predicted routes of 44 molecules are documented in the CAS database, with 36 achieving round-trip scores of 1. Furthermore, the reactions in the predicted routes of 68 molecules are considered correct based on reactions in the CAS database combined with our expertise, with 53 of these molecules attaining round-trip scores of 1. As a result, the predicted routes for 389 of the 501 molecules are considered infeasible, with 185 having round-trip scores of 1 and 204 having scores less than 1. The manual evaluation process requires five days to complete.
We evaluate the ability of synthesizability evaluation metrics to identify two types of routes: feasible and infeasible. For feasible routes, if the round-trip score is 1, it indicates that our forward reaction model successfully simulates the route to synthesize the target molecule, which is considered a successful identification. For infeasible routes, if the round-trip score is smaller than 1, it indicates that the forward reaction model fails to synthesize the target molecule by simulating the route, which is also counted as a successful identification. Based on this, we define the following terms:
True Positive (TP). Correctly identified feasible routes (round-trip score 1 for actual feasible routes): 599.
True Negative (TN). Correctly identified infeasible routes (round-trip score 1 for actual infeasible routes): 204.
False Positive (FP). Incorrectly identified feasible routes (round-trip score 1 for actual infeasible routes): 185.
False Negative (FN). Incorrectly identified infeasible routes (round-trip score 1 for actual feasible routes): 39.
Since the search success rate does not evaluate the feasibility of predicted routes, we define these terms as follows:
True Positive (TP). Feasible routes correctly identified as successful search: 638.
False Positive (FP). Infeasible routes incorrectly identified as successful search (routes generated but are infeasible): 389.
We find that the average SA score is 2.68 for molecules with feasible predicted routes and 2.73 for those with infeasible predicted routes. Since the average SA scores for the two groups are similar, we conclude that the SA score lacks the ability to differentiate between feasible and infeasible routes. Therefore, we do not include it as a baseline.
Additionally, for the search success rate, meaningful TN and FN are absent. Therefore, we use to compare these synthesizability evaluation metrics. Besides, we provide , , and for round-trip score.
| Metric | Accuracy | Precision | Recall | F1 Score |
| Search Success Rate | - | 62.1% | - | - |
| Round-trip Score | 78.2% | 76.4% | 93.9% | 84.2% |
@@ Results.
Table 1 compares the precision of the round-trip score with the search success rate. The results clearly show that the round-trip score surpasses the search success rate, highlighting the forward reaction model’s strength in evaluating route feasibility. This advantage is particularly critical in minimizing false positives, as incorrectly identifying infeasible routes as feasible undermines the reliability of synthesizability evaluation metrics.
Additionally, with a recall of 93.9%, the round-trip score effectively identifies the majority of feasible routes, further demonstrating its reliability in route feasibility assessment. Moreover, as more reaction data becomes available, the accuracy of the forward reaction model is expected to improve, leading to even more reliable round-trip score evaluations.
4.2 Benchmarking Generated Molecules with Round-trip Score
In this section, we utilize the round-trip score to assess the synthesizability of molecules generated by SBDD models.
Settings.
To train a more powerful retrosynthesis model, we split the 107,354 molecules described in Section 4.1 into training, validation, and testing sets, allocating 107,253, 100, and 1 molecule(s) respectively. We employ the same forward reaction prediction model described in Section 4.1. During the search process, we set the beam size to 5 and limit the depth of each route to a maximum of 15. Due to computational constraints, we are unable to use a beam size of 50 as employed in retrosynthesis evaluation (Dai et al., 2019). Besides, our approach generates about 5 synthetic routes per molecule, in contrast to previous methods in retrosynthetic planning that typically produce only one route. This offers a more comprehensive evaluation compared to the search success rate metric used in earlier studies (Chen et al., 2020) for evaluating search algorithms.
Baselines.
For our evaluation, we employ a diverse set of state-of-the-art SBDD models, including LiGAN (Ragoza et al., 2022), AR (Luo et al., 2021), Pocket2Mol (Peng et al., 2022), FLAG (Zhang et al., 2022), TargetDiff (Guan et al., 2023a), DrugGPS (Zhang & Liu, 2023), and DecompDiff (Guan et al., 2023b). These models are trained and tested using the CrossDocked dataset (Francoeur et al., 2020), which comprises an extensive collection of 22.5 million protein-molecule structures. Our experimental setup involves randomly selecting 100,000 protein-ligand pairs from this dataset for training purposes. For testing, we draw 100 proteins from the remaining data points. To ensure a comprehensive evaluation, we randomly sample 100 molecules for each protein pocket in the test dataset, resulting in a total of 10,000 molecules. Additionally, we also verify the validity and plausibility of these molecules. After that, we employ our retrosynthetic planner to generate synthetic routes for them.
Metrics.
We calculate the average number of atoms in the generated molecules and the proportion that are starting materials. Besides, we calculate the percentage of molecules for which at least one of the top-k predicted synthetic routes achieves a round-trip score of 1. Please note that for molecules, which are starting materials, their round-trip scores are set to 1 without predicting any synthetic routes for them.
| Model | Average Number of Atoms | Ratio of Starting Materials | Top-1 | Top-2 | Top-3 | Top-4 | Top-5 |
| LiGAN | 21.17 | 1.66% | 2.46% | 2.60% | 2.73% | 2.79% | 2.87% |
| TargetDiff | 24.46 | 2.05% | 2.81% | 3.13% | 3.30% | 3.35% | 3.41% |
| DecompDiff | 28.34 | 0.53% | 2.84% | 3.46% | 3.78% | 4.01% | 4.05% |
| DrugGPS | 23.36 | 5.54% | 7.49% | 7.90% | 8.23% | 8.35% | 8.41% |
| AR | 17.98 | 4.67% | 7.55% | 8.32% | 8.64% | 8.87% | 9.03% |
| FLAG | 22.42 | 10.35% | 13.40% | 14.36% | 14.73% | 15.01% | 15.22% |
| Pocket2Mol | 18.53 | 14.75% | 19.45% | 20.77% | 21.45% | 21.84% | 22.05% |
Results.
Based on the results presented in Table 2, we can draw several conclusions. There is a significant variation in performance across different SBDD models. As the average number of atoms in the generated molecules increases, the ratio of starting materials and the top-k performance. The top-5 performance ranges from 2.87% for LiGAN to 22.05% for Pocket2Mol, indicating a substantial difference in the models’ abilities to generate synthetically accessible molecules. Pocket2Mol consistently outperforms other models across all metrics, with 22.05% of its generated molecules having at least one synthetic route with a round-trip score of 1 among the top-5 predictions. The improvement in performance from top-1 to top-5 suggests that considering multiple top predictions can significantly increase the likelihood of finding feasible synthetic routes.
Notably, the performance ranking of models remains consistent across all top-k evaluations. The performance increase from top-4 to top-5 is less than 0.25%, indicating that the performance is approaching saturation. Additionally, as shown in Table 4 in Appendix B, the performance gap between models generally widens as increases, with most gaps showing an upward trend. These observations suggest that our chosen beam size of 5 is sufficient to provide an accurate ranking of each model’s performance. This consistency in ranking and the approaching saturation point lend credibility to our evaluation methodology and the reliability of our comparative analysis.
The analysis of molecular properties from various generative models, as presented in Table 5, reveals that superior molecular properties do not always correlate with better synthesizability. Even for the best-performing model, a considerable portion of generated molecules still lack high-quality synthetic routes, indicating room for improvement in generating synthetically accessible molecules in SBDD tasks. These findings underscore the importance of evaluating synthetic accessibility in SBDD models and highlight the potential of using top-k predictions to identify feasible synthetic routes for generated molecules.
@@ 5 Related Work
Synthesizable Drug Design.
Existing methods for synthesizable drug design can be divided into two categories. The first category (Guo & Schwaller, 2024) directly optimizes molecular synthesizability by improving metrics such as the SA score. The second category (Bradshaw et al., 2019b; Vinkers et al., 2003; Gottipati et al., 2020; Horwood & Noutahi, 2020; Gao et al., 2022; Cretu et al., 2024; Luo et al., 2024; Koziarski et al., 2024; Seo et al., 2024) generates the final molecule step-by-step, either from building blocks or existing molecules, using defined synthesis steps and optimizing the process with reward-based methods.
@@
Reaction prediction models are classified into template-based and template-free approaches. Template-based methods (Wei et al., 2016; Segler & Waller, 2017; Qian et al., 2020; Chen & Jung, 2022) predict templates to generate products. Template-free methods are more diverse, using two-stage processes to identify reaction centers and modify bonds (Jin et al., 2017), framing prediction as sequence-to-sequence or graph-to-sequence tasks (Yang et al., 2019; Schwaller et al., 2019a; Tetko et al., 2020; Irwin et al., 2022; Lu & Zhang, 2022; Zhao et al., 2022), or performing direct graph transformations on reactants (Bradshaw et al., 2019a; Do et al., 2019; Sacha et al., 2021; Bi et al., 2021).
@@
Retrosynthesis models are classified into three types: template-free, semi-template-based, and template-based methods, often differentiated by their use of atom mapping information. Template-free methods treat retrosynthesis as a translation (Karpov et al., 2019) or graph edit problem (Sacha et al., 2021). Template-based methods (Segler & Waller, 2017; Dai et al., 2019) utilize atom mapping to build reaction templates, framing retrosynthesis as template classification or retrieval. Semi-template-based methods (Shi et al., 2020; Somnath et al., 2021) often adopt a two-stage approach: identifying reaction centers in the product and breaking it into synthons, then transforming synthons into reactants.
@@
Various search algorithms have been developed for synthetic planning, including beam search, neural A* search (Chen et al., 2020; Han et al., 2022; Xie et al., 2022), Monte Carlo Tree Search (Segler et al., 2018; Hong et al., 2021), and reinforcement learning-based methods (Yu et al., 2022). Other notable contributions (Kishimoto et al., 2019; Heifets & Jurisica, 2012; Kim et al., 2021; Hassen et al., 2022; Li et al., 2023; Zhang et al., 2023; Liu et al., 2023a; Lee et al., 2023; Yuan et al., 2024; Tripp et al., 2024) aim to efficiently navigate the reaction space and prioritize promising synthetic routes.